# Analyzing Probabilistic Logic Shields for Multi-Agent Reinforcement Learning

### v2

@@ 1 Introduction
One recent proposal that uses shielding to represent safety constraints is probabilistic logic shields [PLS, 39] whose semantics is based on probabilistic logic (PL) programming [9]. PLS constrains an agent’s policy to comply with formal specifications probabilistically. The specification of constraints (the shield), is defined within the exploration and learning pipeline. PLS comes with various advantages over other techniques: ($i$) the shields are applied at the policy level instead of individual actions capturing the probabilistic evaluation of safety (in expectation) in contrast to deterministic shields that are based on hard-rejection [16, 17]; ($ii$) PLS is less demanding in terms of requiring knowledge of the underlying MDP compared to previous methods [e.g., 16, 17, 7]; ($iii$) being end-to-end differentiable, it integrates logical semantics with deep learning architectures seamlessly; and ($iv$) it provides safety guarantees within single-agent settings.
@@
    We show that PLS can be used as an equilibrium selection mechanism, a key challenge in MARL, under various game-theoretic settings, such as coordination, spatio-temporal coordination, social dilemmas, cooperation under uncertainty, and mixed-motive problems. We provide empirical evidence across various $n$-player environments including an extensive-form game (Centipede), a stochastic game (Extended Public Goods Game), a simultaneous game (Stag-Hunt), and its grid-world extension (Markov Stag-Hunt). Moreover, we investigate the impact of smaller (weak, less specific) and larger (strong, more specific) shields;
@@ 2 Preliminaries
We start by specifing what we mean by RL safety, provide background knowledge on the game-theoretic environments we use, and finally a quick introduction to PLS.
@@ 2.1 Safety Definition
We expand the interpretation of ‘safety’ in RL beyond that of works like ElSayed-Aly et al. [12] and Yang et al. [39] (i.e. informally, the agent should not perform actions that may lead to something ‘harmful’), and more towards an alternative conception provided in Gu et al. [15, pg. 3], namely, they act, reason, and generalize obeying human desire. We formalize this in the vein of Alshiekh et al. [1]:
@@
Such formal specifications may represent any constraints including arbitrary temporal logic formulae or, in multi-agent settings, social norms or equilibrium conditions, which may not necessarily be ‘harmful’, but may instead ascribe normative or even desirable behavior. In our case, $\mathcal{T}$ is encoded using a PL program.
2.2 Game Theory
A normal-form game (NFG) is a tuple $\langle N,\mathbf{A},\mathbf{u}\rangle$ where $N=\{1,...,n\}$ is a finite set of agents (or players), $\mathbf{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$ is a finite set of action profiles $\mathbf{a}=(a_{1},...a_{n})$ with $\mathcal{A}_{i}$ being the set of player $i$’s actions, and $\mathbf{u}=(u_{1},...,u_{n})$ is a profile of utility functions $u_{i}:\mathbf{A}\rightarrow\mathbb{R}$. A strategy profile $\mathbf{s}=(s_{1},...s_{n})$ is called a Nash equilibrium if for each player $k\in N$, the strategy $s_{k}$ is a best response to the strategies of all the other players $s_{i\in N\setminus\{k\}}$.
@@ 2.3 Probabilistic Logic Shielding
We summarize the approach of Yang et al. [39], known as Probabilistic Logic Shielding (PLS), a method for incorporating probabilistic safety in RL, and refer the reader to the original work for a more detailed treatment. Unlike traditional rejection-based shields that fully block unsafe actions and limit exploration [1], PLS adjusts the probability of unsafe actions proportional to their risk, allowing them to occasionally occur during training. This lets agents learn from unsafe actions, especially under soft constraints.
Assume a model $P$ that assigns probabilities $P(\texttt{safe}|s,a)$ representing the likelihood that taking action $a$ in state $s$ is safe. $P$ may not necessarily represent the whole environment or transition model (i.e. it need not be a full representation of the underlying MDP), but only encode relevant knowledge about the safety of certain states and how they relate to an agent’s current policy $\pi$. The safety of $\pi$ is the sum of the disjoint probabilities of the safety of taking each action:
@@
To obtain a policy that favors safer actions under $P$, we can marginalize actions out w.r.t. $\pi$. This leads us to the notion of probabilistic shielding: formally, given a base policy $\pi$ and a probabilistic safety model $P(\mathtt{safe}|s,a)$, the shielded policy is:
@@
Yang et al. [39] choose to implement the model $P$ in ProbLog [9] for a number of reasons, including that it is easily differentiable and allows easy modeling and planning. This means our safety constraints are to be specified symbolically through PL predicates and relations.
Within PLS, a shield is synonymous with a ProbLog program $\mathcal{T}$ that receives two sets of inputs: the first is the action distribution from a policy $\pi(s)$, and the second is a set of probabilistic fact valuations encoding safety-related information that may come from the agent or the environment – this may differ from the state information used to compute the base policy $\pi$. The appropriate ProbLog program $\mathcal{T}$ induces a probabilistic measure $\mathbf{P}_{\mathcal{T}}$, and we can (conditionally) query the program to give us various information including the action safety (for each action $a$) as $P(\texttt{safe}|s,a)=\mathbf{P}_{\mathcal{T}}(\texttt{safe}|a)$, the safety of the whole policy under $s$ as $P_{\pi}(\texttt{safe}|s)=\mathbf{P}_{\mathcal{T}}(\texttt{safe})$ and the safe/shielded policy $\pi^{+}$ as $P_{\pi}(a|s,\texttt{safe})=\mathbf{P}_{\mathcal{T}}(a|\texttt{safe})$.
@@ 3 Probabilistic Logic Shields for Multi-Agent Reinforcement Learning
We present a general SMARL framework in Section 3.1, applying (probabilistic) shielding to multi-agent algorithms. Next, in Section 3.2, we introduce PLTD, which incorporates logical constraints directly into the TD-learning process. Finally, Section 3.3 outlines the encoding of safety constraints as PL programs.
@@ 3.1 General Framework Description
In Probabilistic Logic SMARL (PL-SMARL), at least one of the RL agents use PLS, optionally sharing policies and/or parameter updates. Two variations (using neural network approximators) of Independent Q-learning (IQL) [31] are introduced as Shielded IQL (SIQL), without parameter sharing, and Shielded Parameter-Sharing IQL (PSQL), sharing all parameters. Likewise, three variants of Independent PPO (IPPO) [10] are examined: Shielded IPPO (SIPPO), Critic-Sharing PPO (CSPPO) and Actor-Critic-Sharing PPO (ACSPPO).
@@ 3.1.1 Safety Guarantees
A proof is provided in the supplementary materials. This is similar to the first case of ElSayed-Aly et al. [12]’s correctness proof of the safety of factored shields. However, unlike factored shields, PLS does not guarantee per-step safety; rather, it ensures that the expected safety under the shielded policy is higher than that of the base policy.
@@ 3.2 PL-Shielded Temporal Difference Learning (PLTD)
Yang et al. [39] define their shielded algorithm, PLPG, as an extension of policy gradient methods, specifically, PPO [27]. Thus, to use it with DQN [21], SARSA [41] (or other TD-learning algorithms), we begin by formalizing how a ProbLog shield is used in conjunction with TD-learning. An important consideration for extending PLS to TD-learning is whether the algorithm is on-policy or off-policy. Yang et al. [39] note that for off-policy algorithms to converge, both the exploration and learned policies must cover the same state-action space – this assumption must be handled with care, particularly in the design of exploration strategies. We propose both on- and off-policy algorithms that incorporate safety constraints into Q-learning.
@@ 3.2.1 Objective Function
PLPG includes a shielded policy gradient $[\nabla_{\theta}\log P_{\pi^{+}}(\mathtt{safe}|s)]$ and a safety gradient penalty $[-\log P_{\pi^{+}}(\mathtt{safe}|s)]$. Since TD-methods do not use policy gradients, we must rely on the latter to introduce information about safety constraints into the objective function.
@@
Under perfect sensor information, PLTD reduces to vanilla TD-learning within the (reachable) set of safe states, and thus in limited settings is guaranteed to converge to an optimal safe policy [33].
@@
A proof of this is in the supplementary material. For deep neural Q-networks, global convergence is currently not known; in that regime we can adopt the usual stabilization tricks (target network, replay buffer, Double-DQN). While we observe empirical stability of PLTD, a complete theoretical analysis remains an open question.
@@ 3.3 Shield Construction using ProbLog
A shield in PLS is defined using a ProbLog program [9], denoted $\mathcal{T}$. Each consists of three components: ($i$) an annotated disjunction $\Pi_{s}$ of predicates whose values are set to a base action distribution $\pi$, ($ii$) a set of constraint-related inputs describing the current state $\mathbf{H}_{s}$, and ($iii$) a set of sentences $\mathcal{KB}$ (for knowledge base) that specify the agent’s constraints, including a definition for a predicate ‘safe_next’ (the predicted safety of the next state under $\pi$). Again, ‘safety’ broadly refers to any probabilistic constraint satisfaction goal (ref. Sec 2.1).
An exemplary shield (Shield 1) for constraining agent behavior towards the mixed Nash equilibrium in Stag-Hunt games is described as such: The probabilistic valuations of the ‘action($\cdot$)’ predicates are inferred from $\pi$. The ‘sensor($\cdot$)’ predicates are set to the normalized absolute difference between the precomputed mixed Nash strategy and the mean of the agent’s historical actions. The predicate ‘safe_next’ is inferred via PL semantics [8] and determines the safety of the next state under $\pi$ – safer states are those where the stated normalized absolute differences are low. For brevity, we describe subsequent shields in-line rather than displaying its full ProbLog source.
@@ 3.3.1 Deterministic Shields
Using PLS, agents may be made to be deterministic in their exploration and deployment which may be useful when training agents in environments requiring hard constraints. This can be formulated as the following (proven in the supplementary):
@@ 4 Results
We organize our experiments around key multi-agent phenomena relevant to several open MARL challenges. Rather than focusing on games individually, each section centers on a particular strategic challenge (e.g., equilibrium selection, social dilemmas) and uses one or more games as illustrative testbeds. For the sake of clarity, details are moved to the supplementary material, such as implementation details, shield programs for each experiment, a validation of PLTD in a single-agent experiment, all hyperparameters and training curves. All displayed results are averaged over 5 random seeds to account for variability in training. All reported errors correspond to standard deviations across runs and respective episodes/steps. Our code can be found at https://github.com/satchitchatterji/ShieldedMARLThesis.
@@ 4.1 Coordination and Equilibrium Selection (Stag-Hunt)
We use Stag-Hunt to test the ability of PLS to guide agents towards normative behaviors in games with multiple Nash equilibria. IPPO serves as the baseline, while SIPPO agents are equipped with a shield constraining the agents towards either the pure ($\mathcal{T}_{pure}$) or mixed Nash equilibrium ($\mathcal{T}_{mixed}$). $\mathcal{T}_{pure}$ simply tells the agent to not play $Hare$, making it a deterministic shield according to Proposition 3.5. $\mathcal{T}_{mixed}$, used as an example in Section 3.3, defines safety as the absolute difference between the precomputed mixed Nash strategy and the mean of the agent’s historical actions.
@@
IPPO 1.99$\pm$0.03 1.99$\pm$0.02 0.01$\pm$0.01
SIPPO ($\mathcal{T}_{pure}$) 4.00$\pm$0.00 4.00$\pm$0.00 1.00$\pm$0.00
SIPPO ($\mathcal{T}_{mixed}$) 2.57$\pm$0.48 2.63$\pm$0.43 0.58$\pm$0.08
Table 1 shows the results this setting. The IPPO agents converge to the non-cooperative Nash equilibrium, consistently playing Hare with a reward of 2 per step. These unshielded agents fail to reach the mixed Nash equilibrium likely due to the instability of PPO’s stochastic policies – small deviations push agents toward pure strategies, often resulting in the non-cooperative equilibrium. In contrast, agents using $\mathcal{T}_{pure}$ quickly and reliably play Stag, earning a higher reward of 4. Agents using $\mathcal{T}_{mixed}$ successfully adopt a mixed strategy, as indicated by the expected rewards and cooperation measures, despite high variability, likely caused by the instability of the normative strategy relative to those under the two pure Nash equilibria.
@@ 4.2 Temporal Coordination (Centipede)
IPPO 42.35$\pm$36.62 42.83$\pm$35.81 0.83$\pm$0.23
SIPPO 100.50$\pm$0.00 100.50$\pm$0.00 1.00$\pm$0.00
IQL ($\epsilon$-greedy) 34.62$\pm$46.59 34.70$\pm$46.53 0.68$\pm$0.23
SIQL ($\epsilon$-greedy) 100.50$\pm$0.00 100.50$\pm$0.00 1.00$\pm$0.00
IQL (softmax) 1.73$\pm$1.01 30.10$\pm$38.61 0.73$\pm$0.21
SIQL (softmax) 100.50$\pm$0.00 100.50$\pm$0.00 1.00$\pm$0.00

### v3

@@ 1 Introduction
One recent proposal that uses shielding to represent safety constraints is probabilistic logic shields [PLS, 46] whose semantics are based on probabilistic logic (PL) programming [10]. PLS constrains an agent’s policy to comply with formal specifications probabilistically. The specification of constraints, (the shield), is defined within the exploration and learning pipeline. PLS offers policy-level evaluation of safety in contrast to action-level hard-rejection shields [e.g., 20, 21, 8], differentiability, modest requirements on knowing the MDP, and safety guarantees within single-agent RL.
@@
    We show that PLS can be used as an equilibrium selection mechanism, a key challenge in MARL, under various game-theoretic settings, such as coordination, spatio-temporal coordination, social dilemmas, cooperation under uncertainty, and mixed-motive problems. We provide strong empirical evidence across various $n$-player environments including an extensive-form game (Centipede), a stochastic game (Extended Public Goods Game), a simultaneous game (Stag-Hunt), and its grid-world extension (Markov Stag-Hunt). Moreover, we investigate the impact of smaller (weak, less specific) and larger (strong, more specific) shields;
@@ 2.1 Safety Definition
We expand the interpretation of ‘safety’ in RL beyond that of works like ElSayed-Aly et al. [13] and Yang et al. [46] (i.e. informally, the agent should not perform actions that may lead to something ‘harmful’), and more towards an alternative conception provided in Gu et al. [17, pg. 3], namely, they act, reason, and generalize obeying human desire. We formalize this in the vein of Alshiekh et al. [2]:
@@
$\mathcal{T}$ may represent any set of constraints, (e.g. arbitrary temporal logic formulae). These may not necessarily be ‘harmful’, but may instead ascribe normative or desirable behavior, such as social norms or equilibrium conditions. In our case, $\mathcal{T}$ is a ProbLog program.
2.2 Multi-Agent Reinforcement Learning
Multi-Agent Reinforcement Learning (MARL) generalizes RL to environments with multiple interacting agents, making the learning problem non-stationary since each agent’s dynamics depend on the evolving policies of others ([1] provides a detailed introduction). Formally, MARL is often modeled as a stochastic game in which each agent $i$ selects actions according to a local policy $\pi_{i}$, and inducing a joint policy $\vec{\pi}=\langle\pi_{1},\dots,\pi_{n}\rangle$ determining the next state and rewards. A common baseline is Independent Q-Learning [IQL, 37], where each agent maintains its own Q-function and updates it independently using local observations and rewards. Similarly, Independent Proximal Policy Optimization [IPPO, 11] adapts PPO to the multi-agent case by training each one with a separate policy-gradient update. In practice, MARL algorithms may employ parameter sharing, where agents use a common neural architecture for policies and/or critics, differing only through their inputs (e.g., observations). This has been shown to improve sample efficiency and facilitates coordination, but may reduce policy diversity and lead to homogenized behaviors [1, 18].
2.3 Game-Theoretic Environments
A normal-form game (NFG) is a tuple $\langle N,\mathbf{A},\mathbf{u}\rangle$ where $N=\{1,...,n\}$ is a finite set of agents (or players), $\mathbf{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$ is a finite set of action profiles $\mathbf{a}=(a_{1},...a_{n})$ with $\mathcal{A}_{i}$ being the set of player $i$’s actions, and $\mathbf{u}=(u_{1},...,u_{n})$ is a profile of utility functions $u_{i}:\mathbf{A}\rightarrow\mathbb{R}$. A strategy profile $\mathbf{s}=(s_{1},...s_{n})$ is called a Nash equilibrium if for each player $k\in N$, the strategy $s_{k}$ is a best response to the strategies of all the other players $s_{i\in N\setminus\{k\}}$ [24, 25].
@@ 2.4 Probabilistic Logic Shielding
We summarize Yang et al. [46]’s approach known as Probabilistic Logic Shielding (PLS), a method for incorporating probabilistic safety in RL, and refer the reader to the original work for a detailed treatment. Unlike earlier rejection-based shields that fully block unsafe actions and limit exploration [2], PLS reduces the probability of unsafe actions proportional to their risk, allowing them to occasionally occur. Under soft constraints, this lets agents learn from unsafe actions.
Assume a model $P$ such that $P(\texttt{safe}|s,a)$ represents the likelihood that taking action $a$ in state $s$ is safe. The safety of $\pi$ is the sum of the disjoint probabilities of the safety of taking each action:
@@
By marginalizing the base policy over the safety model, we obtain a reweighted policy that favors safer actions – this is called probabilistic shielding. Formally, given a base policy $\pi$ and a probabilistic safety model $P$, the shielded policy $\pi^{+}$ is constructed as:
@@
Succinctly, PLS re-normalizes the action distribution to make unsafe actions less likely. Yang et al. [46] chose to implement the model $P$ in ProbLog [10] for a number of reasons, including that it is easily differentiable and allows easy modeling and planning. This means our safety constraints are to be specified symbolically through PL predicates and relations. Within this paper, a shield is thus synonymous with a ProbLog program $\mathcal{T}$ (further details in Section 3.3), which induces a probabilistic measure $\mathbf{P}_{\mathcal{T}}$. We can (conditionally) query this program to give us various information including the action safety (for each action $a$) as $P(\texttt{safe}|s,a)=\mathbf{P}_{\mathcal{T}}(\texttt{safe}|a)$, the safety of the whole policy under $s$ as $P_{\pi}(\texttt{safe}|s)=\mathbf{P}_{\mathcal{T}}(\texttt{safe})$ and the safe/shielded policy $\pi^{+}$ as $P_{\pi}(a|s,\texttt{safe})=\mathbf{P}_{\mathcal{T}}(a|\texttt{safe})$.
@@ 3 Probabilistic Logic Shields for Multi-Agent Reinforcement Learning
We present a general SMARL framework in Section 3.1, applying (probabilistic) shielding to multi-agent algorithms. Next, in Section 3.2, we introduce PLTD, which incorporates logical constraints directly into the TD-learning process. Finally, Section 3.3 outlines the encoding of safety constraints as ProbLog programs.
@@ 3.1 General Framework Description
Though SMARL is agnostic to the underlying MARL algorithm, we experiment with two: IQL and IPPO (introduced in Section 2.2). IQL with parameter sharing will be called Parameter Sharing IQL (PSQL), and IPPO will be called either Critic Sharing IPPO (CSPPO) or Actor-Critic Sharing IPPO (ACSPPO), depending on whether the actors and/or critics are shared. In Probabilistic Logic SMARL (PL-SMARL), at least one of the RL agents uses PLS, optionally sharing policies and parameter updates. PL-SMARL algorithms will be denoted ‘Shielded $X$’ (SXX); e.g., ‘Shielded IQL’, (SIQL) when $X=\text{IQL}$, or ‘Shielded CSPPO’, (SCSPPO) when $X=\text{CSPPO}$.
@@ 3.1.1 Safety Guarantees
A proof is provided in the supplementary materials. This is similar to the first case of ElSayed-Aly et al. [13]’s correctness proof of the safety of factored shields. However, unlike factored shields, PLS does not guarantee per-step safety; rather, it ensures that the shielded algorithm’s safety is higher than that of the base algorithm in expectation.
@@ 3.2 PL-Shielded Temporal Difference Learning (PLTD)
Yang et al. [46] define their shielded algorithm, probabilistic logic policy gradients (PLPG), as an extension of policy gradient methods, specifically, PPO [33]. Thus, to use it with DQN [27], SARSA [48] (or other TD-learning methods), formalize the use of a ProbLog shield in conjunction with TD-learning. An important consideration for extending PLS to TD-learning is whether the algorithm is on-policy or off-policy. Yang et al. [46] note that for off-policy algorithms to converge, the exploration and learned policies must cover the same state-action space – this assumption must be handled with care in the design of exploration strategies. We propose both on- and off-policy algorithms that incorporate safety constraints into Q-learning.
@@ 3.2.1 Objective Function
PLPG includes a shielded policy gradient $\nabla_{\theta}\log P_{\pi^{+}}(\mathtt{safe}|s)]$ and a safety gradient penalty $\nabla_{\theta}\log P_{\pi^{+}}(\mathtt{safe}|s)]$. Since TD-methods do not use policy gradients, we must rely on a modified version of the latter to introduce safety constraints into the objective function.
@@
The quantity $\log P_{\pi^{+}}(\mathtt{safe}|s)]$ is interpreted as the probability that $\pi^{+}$ satisfies the safety constraints [46]. Under perfect sensor information, PLTD reduces to vanilla TD-learning within the reachable set of safe states. We thus get the following convergence property (following Tsitsiklis and Van Roy [39]):
@@
We prove this in the supplementary. For neural Q-networks, global convergence is currently unsolved, though we can adopt common DQN stabilization tricks (summarized in [19]). While we observe empirical stability of PLTD, a complete theoretical analysis is left for future work.
@@ 3.3 Shield Construction using ProbLog
A shield in PLS is defined using a ProbLog program [10], denoted $\mathcal{T}$. Each consists of three components: ($i$) an annotated disjunction $\Pi_{s}$ of predicates whose values are set to a base action distribution $\pi$, ($ii$) a set of constraint-related inputs describing the current state $\mathbf{H}_{s}$, and (iiiiii) a set of sentences $\mathcal{KB}$ (for knowledge base) that specify the agent’s constraints, including a definition for a predicate ‘safe_next’ (the predicted safety of the next state under $\pi$). Again, ‘safety’ broadly refers to any probabilistic constraint satisfaction goal (ref. Sec 2.1).
An exemplary shield (Shield 1) for constraining agent behavior towards the mixed Nash equilibrium in Stag-Hunt games is described as such: The probabilistic valuations of the ‘action($\cdot$)’ predicates are inferred from $\pi$. The ‘sensor($\cdot$)’ predicates are set to the normalized absolute difference between the precomputed mixed Nash strategy and the mean of the agent’s historical actions. The predicate ‘safe_next’ is inferred via ProbLog semantics [9] and determines the safety of the next state under $\pi$ – safer states are those where the stated normalized absolute differences are low. For brevity, we describe subsequent shields in-line rather than displaying its full ProbLog source. For details on how we designed the shields for each experiment, we kindly direct the reader to the supplementary.
@@ 3.3.1 Deterministic Shields
Using PLS, agents may be made to be deterministic which can be useful when training agents in environments requiring hard constraints. This can be formulated as the following (proof in the supplementary):
@@ 4 Results
We organize our experiments around key multi-agent phenomena relevant to several open MARL challenges. Rather than focusing on games individually, each section centers on a particular strategic challenge (e.g., equilibrium selection, social dilemmas) and uses one or more games as illustrative testbeds. For the sake of clarity, details are moved to the supplementary material, such as implementation details, shield programs for each experiment, a validation of PLTD in a single-agent experiment, all hyperparameters and training curves. All displayed results are averaged over 5 random seeds to account for variability in training. All reported errors correspond to standard deviations across runs and respective episodes/steps. Our code is open-sourced at https://github.com/satchitchatterji/ShieldedMARL.
@@ 4.1 Coordination and Equilibrium Selection (Stag-Hunt)
We use Stag-Hunt to test the ability of PLS to guide agents towards normative behaviors in games with multiple Nash equilibria. IPPO serves as the baseline, while SIPPO agents are equipped with a shield constraining the agents towards either the pure cooperative ($\mathcal{T}_{pure}$) or mixed Nash equilibrium ($\mathcal{T}_{mixed}$) – play $Stag/Hare$ 60%/40% of the time with an expected utility of 2.6. $\mathcal{T}_{pure}$ simply tells the agent to not play $Hare$, making it a deterministic shield according to Proposition 3.5. $\mathcal{T}_{mixed}$, used as an example in Section 3.3, defines safety as the absolute difference between the precomputed mixed Nash strategy and the mean of the agent’s historical actions.
@@
[…]
Table 1 shows the results this setting. The IPPO agents converge to the non-cooperative Nash equilibrium, consistently playing Hare with a reward of 2 per step. In contrast, agents using $\mathcal{T}_{pure}$ quickly and reliably play Stag, earning a higher reward of 5. The unshielded agents never converge to the mixed Nash equilibrium, even though PPO is capable of learning stochastic policies – small deviations push these agents toward playing pure strategies, often resulting in the non-cooperative equilibrium. However, shielded PPO agents using $\mathcal{T}_{mixed}$ successfully adopt the mixed strategy, though with high variability due to this attracting nature of the pure equilibria.
@@ 4.2 Temporal Coordination (Centipede)
[…]